OpenStudy (anonymous):

problem is an investor invested a total of $1,500 in two mutual funds. One fund earned a 9% profit while the other earned a 2% profit. If the investor's total profit was $93, how much was invested in the mutual fund? Can someone help me with this please. Not sure how to write this equation out. Thanks

OpenStudy (anonymous):

\[.09x+.02(1500-x)=93\] solve for \(x\)

OpenStudy (anonymous):

reasoning that if you put \(x\) equal the amount invested at 9% then since the total is 1500 the amount invested at 2% must be \(1500-x\)

OpenStudy (anonymous):

and since you know the total is 93, set the expression for the amount of interest equal to 93 and solve probably easier to multiply all by 100 and solve \[9+2(1500-x)=9300\] so you don't have to work with the decimals

OpenStudy (anonymous):

sorry last line should have been \[9x+2(1500-x)=9300\]

OpenStudy (anonymous):

you got it from there? i get \(x=900\)

OpenStudy (anonymous):

so how would the equation be written out. just 9x+ 2(1500-x) = 9300 9x + 3000 - 2x = 9300

OpenStudy (anonymous):

yes, three more steps and you are done

OpenStudy (anonymous):

-7x = 6300 x=900

OpenStudy (anonymous):

\[7x+3000=9300\] \[7x=6300\] \[x=900\]

OpenStudy (anonymous):

but i need the amount the 9% was and for the 2% as well.

OpenStudy (anonymous):

that is why you need to write what \(x\) represents at the start. i put \(x\)= amount invested at 9%

OpenStudy (anonymous):

once we solve for \(x\) we know that the amount invested at 9% was $900 and since the total was $1500 the amount invested at 2% has to be $600

OpenStudy (anonymous):

ok. I get it.. thank you